OpenStudy (anonymous):

Having a problem with a word problem: a rectangular garden has dimensions of 25 feet by 13 feet. A gravel path of consistent width is to be built around the garden. How wide can the path be if there is enough gravel for 480 feet?

OpenStudy (anonymous):

let x =width of the gravel path Draw rough diagram of this labeling the garden as 25 by 13and the width of the gravel path as x You will see the overall dimensions will be (2x+25) by (2x+13) Overall area - garden area = gravel path area (required to be 480 feet (2x+25)(2x+13)-25.13=480 4x^2+26x+50x+325-325=480 4x^2+26x-480=0 Simplify, divide equation by 2: 2x^2+13x-240=0

OpenStudy (anonymous):

middle term the equation

OpenStudy (anonymous):

That is where I end up getting stuck..divide all by 2 then factor?